Earl Ezekiel wants to retire in San Diego when he is 65 years old. Earl is now 50. He believes he will need $300,000 to retire comfortably. To date, Earl has set aside no retirement money. Assume Earl gets 6% interest compounded semiannually. How much must Earl invest today to meet his $300,000 goal?
Answer by Boreal(15235)

P=Po(1+.06/2)^30, the 30 being semiannual compoundings over 15 years.
300000=Po(1.03)^30
300000=Po*2.427. Without rounding divide 300000 by 2.427.
$123,596.03